### v2

@@ 3.4.3 Semantic Segmentation
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=6.5em,font=,, where level=2text width=9.5em,font=,, where level=3text width=8.0em,font=,, where level=4text width=7.5em,font=,, [ Application, ver, color=carminepink!100, fill=carminepink!15, text=black [ Image (§4.1), color=lighttealblue!100, fill=lighttealblue!100, text=black, text width=5em [ Image Classification
4.1.1), color=lighttealblue!100, fill=lighttealblue!60, text=black, text width=8.5em [ Natural Image, color=lighttealblue!100, fill=lighttealblue!40, text=black, text width=6.5em [ e.g. Mamba-ND (Li et al, 2024d) , cause_leaf, text width=28.5em ] ] [ Medical Image, color=lighttealblue!100, fill=lighttealblue!40, text=black, text width=6.5em [ e.g. MamMIL (Fang et al, 2024), MambaMIL (Yang et al, 2024c),
      Vim4Path (Nasiri-Sarvi et al, 2024), CMViM (Yang et al, 2024b) , cause_leaf, text width=28.5em ] ] [ Remote Sensing
Image, color=lighttealblue!100, fill=lighttealblue!40, text=black, text width=6.5em [ e.g. RSMamba (Chen et al, 2024d) , cause_leaf, text width=28.5em ] ] ] [ Image Segmentation
4.1.2), color=lighttealblue!100, fill=lighttealblue!60, text=black, text width=8.5em [ Natural Image, color=lighttealblue!100, fill=lighttealblue!40, text=black, , text width=6.5em [ e.g. MiM-ISTD (Chen et al, 2024e) , cause_leaf, text width=28.5em ] ] [ Medical Image, color=lighttealblue!100, fill=lighttealblue!40, text=black, , text width=6.5em [ e.g. U-Mamba (Ma et al, 2024d), SegMamba (Xing et al, 2024),
      Mamba-UNet (Wang et al, 2024j), Swin-UMamba (Liu et al, 2024c),
      VM-UNet (Ruan and Xiang, 2024), nnMamba (Gong et al, 2024),
      H-vmunet (Wu et al, 2024a), LightM-UNet (Liao et al, 2024) , cause_leaf, text width=28.5em ] ] [ Remote Sensing
Image, color=lighttealblue!100, fill=lighttealblue!40, text=black, text width=6.5em [ e.g. Samba (Zhu et al, 2024b), SpectralMamba (Yao et al, 2024),
      RS3Mamba (Ma et al, 2024e), ChangeMamba (Chen et al, 2024c) , cause_leaf, text width=28.5em ] ] ] [ Object Detection
4.1.3), color=lighttealblue!100, fill=lighttealblue!60, text=black, text width=8.5em [ e.g. FER-YOLO-Mamba (Ma et al, 2024a), SOAR (Verma et al, 2024) , cause_leaf, text width=28.5em ] ] [ Image Generation
4.1.4), color=lighttealblue!100, fill=lighttealblue!60, text=black, text width=8.5em [ e.g. ZigMa (Hu et al, 2024), Gamba (Shen et al, 2024),
      MD-Dose (Fu et al, 2024b), VM-DDPM (Ju and Zhou, 2024) , cause_leaf, text width=28.5em ] ] [ Image Restoration
4.1.5), color=lighttealblue!100, fill=lighttealblue!60, text=black, text width=8.5em [ Natural Image, color=lighttealblue!100, fill=lighttealblue!40, text=black, text width=6.5em [ e.g. UVMNet (Zheng and Wu, 2024), FreqMamba (Zhen et al, 2024),
      MambaUIE (Chen and Ge, 2024), Retinexmamba (Bai et al, 2024) , cause_leaf, text width=28.5em ] ] [ Medical Image, color=lighttealblue!100, fill=lighttealblue!40, text=black, text width=6.5em [ e.g. FDVMNet (Zheng and Zhang, 2024), MSDR (Li et al, 2024i) , cause_leaf, text width=28.5em ] ] [ Remote Sensing
Image, color=lighttealblue!100, fill=lighttealblue!40, text=black, text width=6.5em [ e.g. HSIDMamba (Liu et al, 2024h), SSUMamba (Fu et al, 2024a),
      FMSR (Xiao et al, 2024), GMSR (Wang et al, 2024e) , cause_leaf, text width=28.5em ] ] ] [ Others
4.1.6), color=lighttealblue!100, fill=lighttealblue!60, text=black, text width=8.5em [ e.g. MambaAD (He et al, 2024a), MambaVC (Qin et al, 2024) , cause_leaf, text width=28.5em ] ] ] [ Video (§4.2), color=lightplum!100, fill=lightplum!100, text=black, text width=5em [ Video Understanding
4.2.1), color=lightplum!100, fill=lightplum!60, text=black, text width=8.5em [ e.g. Vivim (Yang et al, 2024e), VideoMamba (Li et al, 2024c),
      VideoMamba Suite (Chen et al, 2024a) , detect_leaf, text width=27.5em ] ] [ Video Generation
4.2.2), color=lightplum!100, fill=lightplum!60, text=black, text width=8.5em [ e.g. ZigMa (Hu et al, 2024), Matten (Gao et al, 2024) , detect_leaf, text width=27.5em ] ] ] [ Point Cloud
4.3), color=lightblue!100, fill=lightblue!100, text=black, text width=5em [ e.g. PointMamba (Liang et al, 2024), PCM (Zhang et al, 2024d),
      3DMambaIPF (Zhou et al, 2024b), 3DMambaC (Li et al, 2024g) , point_leaf, text width=30em ] ] [ Multi-modal
@@
4.4), color=lightgreen!100, fill=lightgreen!100, text=black, text width=5em [ Homogeneous
Multi-modal (§4.4.1), color=lightgreen!100, fill=lightgreen!60, text=black, text width=8.5em [ e.g. MambaDFuse (Li et al, 2024h), FusionMamba (Xie et al, 2024b),
      S4Fusion (Ma et al, 2024b), Fusion-Mamba (Dong et al, 2024),
      CFMW (Li et al, 2024b), Sigma (Wan et al, 2024),
      MambaMorph (Guo et al, 2024b), VMambaMorph (Wang et al, 2024i) , modal_leaf, text width=30em ] ] [ Heterogeneous
Multi-modal (§4.4.2), color=lightgreen!100, fill=lightgreen!60, text=black, text width=8.5em [ e.g. InstructGIE (Meng et al, 2024), VLMamba (Qiao et al, 2024),
      Meteor (Lee et al, 2024), Cobra (Zhao et al, 2024a),
      ReMamber (Yang et al, 2024d), CLIP-Mamba (Huang et al, 2024f),
      SurvMamba (Chen et al, 2024g), RSCaMa (Liu et al, 2024a) , modal_leaf, text width=30em ] ] ] ]
@@
Figure 6: The main content flow and categorization of the application in this survey
4 Application
@@
In this section, we systematically categorize and discuss the diverse vision applications of Mamba. The categorization scheme, along with an overview of the reviewd literature in this survey, is presented in Fig. 6.
@@
Table 4: The Mamba-based methods in image-level vision tasks (Part I). The abbreviations of Block here are VSS: Vision State Space Module, Vim: Vision Mamba, *: modification, FT: Fourier Transform. The abbreviations of Data here are CBCT: Cone Beam Computed Tomography, CXR: Chest X-Ray, CT: Computed Tomography, Derm: Dermatoscope, Echo: Echocardiograms, Endo: Endoscopic Images, FMRI: Fast Magnetic Resonance Imaging, Micro: Microscopy Images, MRI: Magnetic Resonance Imaging, OCT: Optical Coherence Tomography, Path: Pathology, PET: Positron Emission Tomography, Skin: Skin Lesion Images, Spleen: Spleen Images, SVCT: Sparse-view Computed Tomography, UI: Ultrasound Images, ODT: Optical Doppler Tomography, HSI: Hyperspectral Images. The abbreviations of Task here are CLS: Classification, DET: Detection, SEG: Segmentation
Methods Block Data Task Scan Continuity Code
@@
InsectMamba (Wang et al, 2024d) Mamba* Natural Images Insect Pests CLS SD (H) Raster
MamMIL (Fang et al, 2024) Vim* Whole Slide Images Cancer Subtyping BD (H) Raster
MambaMIL (Yang et al, 2024c) Mamba* Whole Slide Images Cancer Subtyping/Survival Prediction SD (H) + Reordering Raster
Vim4Path (Nasiri-Sarvi et al, 2024) Vim Whole Slide Images Cancer Subtyping BD (H) Raster
CMViM (Yang et al, 2024b) Vim 3D Medical Images (MRI & PET) 3D Medical Image CLS BD (H) Raster
nnMamba (Gong et al, 2024) Mamba* 3D Medical Images (CT/MRI) 3D Medical Image CLS BD (H) + Channel Raster
MedMamba (Yue and Li, 2024) VSS* 2D Medical Images (Skin/UI/CXR/Endo/ 2D Medical Image CLS BD (H/V) Raster
OCT/CT/Micro/Path/Derm)
BI-Mamba (Yang et al, 2024f) Vim 2D Medical Images (X-ray) Risk Prediction BD (H) Raster
RSMamba (Chen et al, 2024d) Mamba Remote Sensing Images Remote Sensing Images CLS BD (H/V) + Shuffle Raster
@@
MiM-ISTD (Chen et al, 2024e) VSS* Infrared Images Infrared Image SEG BD (H/W) Raster
U-Mamba (Ma et al, 2024d) Mamba* 2D/3D Medical Images (Endo/Micro/CT/MRI) 2D/3D Medical Image SEG SD (H) Raster
SegMamba (Xing et al, 2024) Mamba* 3D Medical Images (CT/MRI) 3D Medical Image SEG BD (H) + SD (D) Raster
VM-UNet (Ruan and Xiang, 2024) VSS 2D Medical Images (Skin/CT) 2D Medical Image SEG BD (H/V) Raster
Swin-UMamba (Liu et al, 2024c) VSS 2D Medical Images (Endo/Micro/MRI) 2D Medical Image SEG BD (H/V) Raster
nnMamba (Gong et al, 2024) Mamba* 3D Medical Images (CT/MRI) 3D Medical Images SEG/Landmark DET BD (H) + Channel Raster
Mamba-UNet (Wang et al, 2024j) VSS 2D Medical Images (CT/MRI) 2D Medical Image SEG BD (H/V) Raster
Mamba-ND (Li et al, 2024d) Mamba* 3D Medical Images (CT) 3D Medical Images SEG BD (H/V/D) Raster
Semi-Mamba-UNet (Ma and Wang, 2024) VSS 2D Medical Images (MRI) 2D Medical Image SEG BD (H/V) Raster
P-Mamba (Ye and Chen, 2024) Vim 2D Medical Images (Echo) 2D Medical Image SEG BD (H) Raster
Weak-Mamba-UNet (Wang and Ma, 2024) VSS 2D Medical Images (MRI) 2D Medical Image SEG BD (H/V) Raster
LightM-UNet (Liao et al, 2024) VSS 2D/3D Medical Images (CXR/CT) 2D/3D Medical Image SEG BD (H/W) Raster
LW-Mamba-UNet (Wang et al, 2024c) Vim* 2D/3D Medical Images (CT/MRI) 2D/3D Medical Image SEG BD (H) Raster
VM-UNet-V2 (Zhang et al, 2024c) VSS 2D Medical Images (Skin/Endo) 2D Medical Image SEG BD (H/V) Raster
H-vmunet (Wu et al, 2024a) VSS* 2D Medical Images (Skin/Spleen/Endo) 2D Medical Image SEG BD (H/V) Raster
ProMamba (Xie et al, 2024a) Vim 2D Medical Images (Endo) 2D Medical Image SEG BD (H) Raster
TM-UNet (Tang et al, 2024a) VSS* 2D Medical Images (Endo/Micro/CT/MRI) 2D Medical Image SEG BD (H/V) Raster
Mamba-HUNet (Sanjid et al, 2024a) VSS 2D Medical Images (MRI) 2D Medical Image SEG BD (H/V) Raster
UltraLight VM-UNet (Wu et al, 2024b) VSS* 2D Medical Images (Skin) 2D Medical Image SEG BD (H/V) Raster
T-Mamba (Hao et al, 2024) Vim* 3D Medical Images (CBCT) 3D Medical Image SEG BD (H) + FT Raster
@@
ViM-UNet (Archit and Pape, 2024) Vim 2D Medical Images (Cell/Neurite) 2D Medical Images SEG BD (H) Raster
nnU-Net Revisited (Isensee et al, 2024) Mamba* 3D Medical Images (CT/MRI) 3D Medical Images SEG SD (H) Raster
Mamba-AHNet (Sanjid et al, 2024b) VSS* 2D Medical Images (CT) 2D Medical Images SEG BD (H/V) Raster
AC-MambaSeg (Nguyen et al, 2024) VSS* 2D Medical Images (Skin) 2D Medical Images SEG BD (H/V) Raster
HC-Mamba (Xu, 2024) VSS* 2D Medical Images (Skin/CT) 2D Medical Images SEG BD (H/V) Raster
MUCM-Net (Yuan et al, 2024) Mamba 2D Medical Images (Skin) 2D Medical Image SEG SD (H) Raster
@@
UU-Mamba (Tsai et al, 2024) Mamba 2D Medical Images (MRI) 2D Medical Image SEG SD (H) Raster
@@
TokenUnify (Chen et al, 2024f) Mamba 3D Medical Images (Micro) 3D Medical Image SEG SD (H) Raster
Samba (Zhu et al, 2024b) Mamba* Remote Sensing Images Semantic SEG SD (H) Raster
@@
RS3Mamba (Ma et al, 2024e) VSS Remote Sensing Images Semantic SEG BD (H/V) Raster
SpectralMamba (Yao et al, 2024) Mamba* Remote Sensing Images Semantic SEG SD (H) Raster
S2Mamba (Wang et al, 2024b) Mamba* Remote Sensing Images Semantic SEG BD (H/V) Raster
SS-Mamba (Huang et al, 2024d) Mamba Remote Sensing Images Semantic SEG SD (H) Raster
Experimental Mamba (Zhu et al, 2024c) VSS* Remote Sensing Images Semantic SEG BD (H/V/LD/RD) Raster/Zigzag
CM-UNet (Liu et al, 2024e) VSS* Remote Sensing Images Semantic SEG BD (H/V) Raster
RS-Mamba (Zhao et al, 2024b) Mamba* Remote Sensing Images Semantic SEG/Change DET BD (H/V/LD/RD) Raster
@@
ChangeMamba (Chen et al, 2024c) VSS Remote Sensing Images Change DET/Building Damage Assessment BD (H/V) Raster
Mamba-in-Mamba (Zhou et al, 2024c) VSS* Remote Sensing Images Remote Sensing Images CLS H/V/H*/V* Zigzag
3DSS-Mamba (He et al, 2024c) Mamba* Remote Sensing Images Remote Sensing Images CLS H/V Raster
4.1 Image
4.1.1 Classification
In addition to the backbones (Liu et al, 2024g; Zhu et al, 2024a) conducting image classification for representation learning, the scalability of Mamba-based architectures to long sequences has led to their adoption in the analysis of high-dimensional images (e.g., Whole Slide Images, 3D Medical Images and Remote Sensing Images). For instance, MamMIL (Fang et al, 2024) and MambaMIL (Yang et al, 2024c) utilized Mamba to enhance Multiple Instance Learning (MIL) for histopathology image analysis. MamMIL employed a bidirectional state space model similar to Vim and a 2D context-aware block (Shao et al, 2021) to enhance instance aggregation. MambaMIL introduced a new approach named Sequence Reordering Mamba (SR-Mamba), which is sensitive to the order and distribution of instances and leverages the valuable information embedded within these long sequences. Differently, Vim4Path (Nasiri-Sarvi et al, 2024) utilized Vim architecture within the DINO (Caron et al, 2021) framework for representation learning. CMViM (Yang et al, 2024b) proposed a contrastive masked Vim autoencoder tailored for 3D multi-modal data to facilitate Alzheimer’s disease (AD) recognition. Other medical image classification methods incorporated the Mamba architecture with convolutions (Gong et al, 2024; Yue and Li, 2024), channel scanning techniques (Gong et al, 2024), and multi-view fusion (Yang et al, 2024f).
In the remote sensing scenario, RSMamba (Chen et al, 2024d) introduced a position-sensitive dynamic multi-path activation mechanism, encompassing forward, backward, and shuffle paths, to extract precise semantic cues for accurate scene discrimination.
InsectMamba (Wang et al, 2024d) integrated SSMs with convolutions, multi-head self-attention mechanisms, and multilayer perceptions to synergistically leverage the strengths of each encoding strategy for insect pest classification.
4.1.2 Segmentation
Segmentation remains a vital and prominent area in computer vision, holding immense value for diverse real-world applications. Mamba is poised to enhance segmentation tasks due to its capabilities in managing extensive visual data.
MIM-ISTD (Chen et al, 2024e) introduced a novel Mamba-in-Mamba structure, where it employs Outer Mamba to explore the global information and utilizes Inner Mamba to further explore the local information within each visual patch.
Recent advances based on Mamba have appeared in the medical domain, encompassing diverse 2D and 3D medical image datasets Tsai et al (2024); Chen et al (2024f). Several methods (Liu et al, 2024c; Ruan and Xiang, 2024; Ma and Wang, 2024; Wang et al, 2024j; Sanjid et al, 2024a) directly replace CNN-blocks in the U-Net architecture with Mamba-like blocks (Gu and Dao, 2023; Liu et al, 2024g; Zhu et al, 2024a). U-Mamba (Ma et al, 2024d) is the first effective attempt to apply Mamba in medical image segmentation. It introduced a hybrid CNN-SSM block. SegMamba (Xing et al, 2024) introduced the Gated Spatial Convolution and Tri-orientated Spatial Mamba, which utilized an inter-slice path for 3D data. VM-UNet (Ruan and Xiang, 2024), VM-UNet-V2 Zhang et al (2024c), Swin-UMamba (Liu et al, 2024c), Mamba-UNet (Wang et al, 2024j), Mamba-HUNet (Sanjid et al, 2024a), Mamba-Ahnet (Sanjid et al, 2024b), Ac-mambaseg (Nguyen et al, 2024) and HC-Mamba (Xu, 2024) employed VSS as a basic block to capture extensive contextual information. Mamba-UNet (Wang et al, 2024j) was further enhanced through the incorporation of semi-supervised (Ma and Wang, 2024) and weak-supervised settings (Wang and Ma, 2024). P-Mamba (Ye and Chen, 2024), ViM-UNet (Archit and Pape, 2024), LW-Mamba-UNet (Wang et al, 2024c), Pro-Mamba (Xie et al, 2024a), T-Mamba (Hao et al, 2024), TM-UNet (Tang et al, 2024a), UltraLight VM-UNet Wu et al (2024b), and MUCM-Net Yuan et al (2024) employed Mamba-like architecture as a basic block with additional designs for enhancement. Meanwhile, some of these Mamba-based methods (Ma et al, 2024d; Ruan and Xiang, 2024; Liu et al, 2024c; Liao et al, 2024; Wang et al, 2024c; Wu et al, 2024a) conducted extensive experiments on both 2D and 3D medical datasets, which demonstrates the efficacy of Mamba architecture for both 2D and 3D medical images. However, (Isensee et al, 2024) noticed validation pitfalls in medical image segmentation. To avoid these pitfalls, a thorough and comprehensive benchmarking of segmentation methods, including CNN-based, Transformer-based, and Mamba-based approaches, was conducted. The results reveal that most methods introduced in recent years fail to surpass the original nnU-Net baseline (Isensee et al, 2018). This motivates us to think about the principle of Mamba compared to CNN and Transformer, and pay more attention to stringent validation standards.
In the field of remote sensing images, Mamba-based approaches typically leverage its efficiency for processing high-resolution images (Zhu et al, 2024b; Ma et al, 2024e; Liu et al, 2024e) and exploit the spatial-spectral characteristics of images to perform scanning (Wang et al, 2024b; Huang et al, 2024d; He et al, 2024c). SpectralMamba (Yao et al, 2024) proposed gated spatial-spectral merging and piece-wise sequential scanning. Mamba-in-Mamba (Zhou et al, 2024c) proposed a centralized Mamba Cross-Scan (MCS) mechanism with a tokenized Mamba to enhance representative feature generation and concentration. ChangeMamba (Chen et al, 2024c) adopted the VSS architecture as the encoder and spatial-temporal relationship modeling mechanisms to achieve spatial-temporal interaction of multi-temporal features. RS-Mamba (Zhao et al, 2024b) introduced an Omni-directional Selective Scan Module based on Mamba. Regarding various scanning strategies adopted in these works, (Zhu et al, 2024c) conducted a comparative study to evaluate the impact of 22 scanning strategies in Mamba-based methods on semantic segmentation of remote sensing images. The findings indicate that these strategies do not significantly enhance performance in this task, raising questions about the effectiveness of multi-directional scanning approaches.
Table 5: The Mamba-based methods in image-level vision tasks (Part II)
Methods Block Data Task Scan Continuity Code
Object Detection
FER-YOLO-Mamba (Ma et al, 2024a) VSS* Facial Images Facial Expression DET BD (H/V/LD/RD) Raster
VMambaCC (Ma et al, 2024c) VSS Natural Images Crowd Counting BD (H/V) Raster
SOAR (Verma et al, 2024) Vim Remote Sensing Images Small Object DET BD (H) Raster

### v3

@@ 3.3.3. Semantic Segmentation
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=6.5em,font=,, where level=2text width=9.5em,font=,, where level=3text width=8.0em,font=,, where level=4text width=7.5em,font=,, [ Application, ver, color=carminepink!100, fill=carminepink!15, text=black [ Image (§4.1), color=lighttealblue!100, fill=lighttealblue!100, text=black, text width=5.5em [ Image Classification (§4.1.1), color=lighttealblue!100, fill=lighttealblue!60, text=black, text width=10em [ e.g. MambaTSR (Ge et al., 2024), MambaMIL (Yang et al., 2024d), Vim4Path (Nasiri-Sarvi et al., 2024), BI-Mamba (Yang et al., 2024f), MSSM (Li et al., 2024j), RSMamba (Chen et al., 2024a), color=lighttealblue!100, fill=lighttealblue!40, text=black, text width=34em ] ] [ Image Segmentation (§4.1.2), color=lighttealblue!100, fill=lighttealblue!60, text=black, text width=10em [ e.g. HMNet (Xu et al., 2024b), ViM-UNet (Archit and Pape, 2024), Swin-UMamba (Liu et al., 2024g), VM-UNet-V2 (Zhang et al., 2024h), MM-UNet (Du et al., 2024b), Polyp-Mamba (Xu et al., 2024c),
GGVMamba (Zhou et al., 2024d), LKM-UNet (Wang et al., 2024b), SegMamba (Xing et al., 2024), EM-Net (Chang et al., 2024), CAMS-Net (Khan et al., 2025), TP-Mamba (Wang et al., 2024g), PathMamba (Fan et al., 2024),
ShapeMamba-EM (Shi et al., 2024c), Semi-Mamba-UNet (Ma and Wang, 2024), RS3Mamba (Ma et al., 2024d), ChangeMamba (Chen et al., 2024e), MambaHSI (Li et al., 2024e), color=lighttealblue!100, fill=lighttealblue!40, text=black, , text width=34em ] ] [ Object Detection (§4.1.3), color=lighttealblue!100, fill=lighttealblue!60, text=black, text width=10em [ e.g. FER-YOLO-Mamba (Ma et al., 2024a), VMambaCC (Ma et al., 2024e), SOAR (Verma et al., 2024), color=lighttealblue!100, fill=lighttealblue!40, text=black, , text width=34em ] ] [ Image Generation (§4.1.4), color=lighttealblue!100, fill=lighttealblue!60, text=black, text width=10em [ e.g. ZigMa (Hu et al., 2024), VSSM (Lambrechts et al., 2024), color=lighttealblue!100, fill=lighttealblue!40, text=black, , text width=34em ] ] [ Image Restoration (§4.1.5), color=lighttealblue!100, fill=lighttealblue!60, text=black, text width=10em [ e.g. FreqMamba (Zhen et al., 2024), MambaIR (Guo et al., 2024a), DVMSR (Lei et al., 2024), MambaLLIE (Weng et al., 2024), ECMamba (Dong et al., 2024c), Deform-Mamba (Ji et al., 2024),
SSUMamba (Fu et al., 2024c), FMSR (Xiao et al., 2024c), MambaFormerSR (Zhi et al., 2024), VMambaSCI (Zhang et al., 2024c), color=lighttealblue!100, fill=lighttealblue!40, text=black, , text width=34em ] ] [ Others (§4.1.6), color=lighttealblue!100, fill=lighttealblue!60, text=black, text width=10em [ e.g. MambaAD (He et al., 2024a), MTMamba (Lin et al., 2024), Hamba (Dong et al., 2024a), SUM (Hosseini et al., 2025), color=lighttealblue!100, fill=lighttealblue!40, text=black, , text width=34em ] ] ] [ Video (§4.2), color=lightplum!100, fill=lightplum!100, text=black, text width=5.5em [ Video Understanding (§4.2.1), color=lightplum!100, fill=lightplum!60, text=black, text width=10em [ e.g. VideoMamba (Li et al., 2024c), VideoMamba (Park et al., 2024a), MambaTrack (Xiao et al., 2024a), DLT (Jin et al., 2024), VMRNN (Tang et al., 2024b) , color=lightplum!100, fill=lightplum!40, text=black, , text width=34em ] ] [ Video Generation (§4.2.2), color=lightplum!100, fill=lightplum!60, text=black, text width=10em [ e.g. RainMamba (Wu et al., 2024c), SSMDiff (Oshima et al., 2024), ZigMa (Hu et al., 2024), VFIMamba (Zhang et al., 2024e), MMD (Dong et al., 2024b) , color=lightplum!100, fill=lightplum!40, text=black, text width=34em ] ] ] [ Point Cloud (§4.3), color=lightblue!100, fill=lightblue!100, text=black, text width=5.5em [ e.g. PointMamba (Liang et al., 2024), Mamba3D (Han et al., 2024), 3DMambaIPF (Zhou et al., 2024c), LCM (Li et al., 2024i), VoxelMamba (Zhang et al., 2024b), MambaMOS (Zeng et al., 2024) , color=lightblue!100, fill=lightblue!40, text=black, text width=34em ] ] [ Multi-modal (§4.4), color=lightgreen!100, fill=lightgreen!100, text=black, text width=5.5em [ Homogeneous Multi-modal (§4.4.1), color=lightgreen!100, fill=lightgreen!60, text=black, text width=10em [ e.g. MGMF (Wang et al., 2024k), Sigma (Wan et al., 2025) , color=lightgreen!100, fill=lightgreen!40, text=black, text width=34em ] ] [ Heterogeneous Multi-modal (§4.4.2), color=lightgreen!100, fill=lightgreen!60, text=black, text width=10em [ e.g. InstructGIE (Meng et al., 2024), Meteor (Lee et al., 2024), RoboMamba (Liu et al., 2024d), ReMamber (Yang et al., 2024c), RSCaMa (Liu et al., 2024a), CoupledMamba (Li et al., 2024k),
@@
MotionMamba (Zhang et al., 2024f), MambaTalk (Xu et al., 2024a), MambaGesture (Fu et al., 2024b) , color=lightgreen!100, fill=lightgreen!40, text=black, text width=34em ] ] ] ]
@@
Figure 6. The main content flow and categorization of the application in this survey.
4. Applications
@@
In this section, we systematically categorize and discuss the diverse vision applications of Mamba. The categorization scheme, along with an overview of the reviewed literature in this survey, is presented in Fig. 6.
4.1. Image
The processing approaches used in image applications are built upon the techniques employed in visual Mamba backbone networks, including tokenization methods, scanning strategies, and block designs, among others.
4.1.1. Classification
In addition to backbones like Vim (Zhu et al., 2024c) and VMamba (Liu et al., 2024e), which are primarily used for image classification in representation learning, the efficiency and scalability of Mamba-based architectures in handling long sequences have enabled their application across a broad range of image classification tasks (Wang et al., 2024i; Ge et al., 2024; Fang et al., 2024; Yang et al., 2024d; Nasiri-Sarvi et al., 2024; Yang et al., 2024b; Gong et al., 2024; Yue and Li, 2024; Yang et al., 2024f; Li et al., 2024j; Chen et al., 2024a). This makes them particularly effective for high-dimensional image analysis, (e.g., Whole Slide Images, 3D Medical Images, and Remote Sensing Images).
MambaTSR (Ge et al., 2024) leveraged EfficientVMamba (Pei et al., 2024) blocks to achieve efficient and accurate traffic sign recognition.
With regards to high-dimensional image analysis, MambaMIL (Yang et al., 2024d) utilized Mamba to enhance Multiple Instance Learning (MIL) for histopathology image analysis. It introduced a new approach named Sequence Reordering Mamba (SR-Mamba), which is sensitive to the order and distribution of instances and leverages the valuable information embedded within these long sequences. Differently, Vim4Path (Nasiri-Sarvi et al., 2024) utilized Vim architecture within the DINO (Caron et al., 2021) framework for representation learning. BI-Mamba (Yang et al., 2024f) employed parallel forward and backward Mamba blocks to effectively capture long-range dependencies in multi-view high-resolution chest X-rays. MSSM (Li et al., 2024j) introduced a multi-scale SSM module for 4D rs-fMRI data, specifically applied to the classification of major depressive disorder.
In the remote sensing scenario, RSMamba (Chen et al., 2024a) introduced a position-sensitive dynamic multi-path activation mechanism, encompassing forward, backward, and shuffle paths, to extract precise semantic cues for accurate scene discrimination.
@@ 4.1.1. Classification
Table 6. The Mamba-based methods in image modality vision tasks. The abbreviations of Data here are CXR: Chest X-Ray, CT: Computed Tomography, MRI: Magnetic Resonance Imaging, rs-fMRI: Resting-state Functional Magnetic Resonance Imaging, EM: Electron Microscopy, Echo: Echocardiograms, Micro: Microscopy Images, Skin: Skin Lesion Images, Path: Pathology. The abbreviations of Task here are CLS: Classification, DET: Detection, SEG: Segmentation, SU: Scene Understanding. Tok: Tokenization.
Methods Data Task Tok. Scan Strategy Venue Code
Direction Axis Continuity Sampling
@@
MambaTSR (Ge et al., 2024) Natural Images Traffic Sign CLS 2D BD H/V Raster Global Neurocomputing24
MambaMIL (Yang et al., 2024d) Whole Slide Images Cancer Subtyping/Survival Prediction 1D SD H Raster + Reordering Global MICCAI24
Vim4Path (Nasiri-Sarvi et al., 2024) Whole Slide Images Cancer Subtyping 1D BD H Raster Global CVPR24 Workshop
BI-Mamba (Yang et al., 2024f) 2D Medical Images (X-ray) Risk Prediction 1D BD H Raster Global MICCAI24 Oral
MSSM (Li et al., 2024j) 4D Medical Images (rs-fMRI) Major Depressive Disorder CLS 1D SD H Raster Global KDD24 Workshop
RSMamba (Chen et al., 2024a) Remote Sensing Images Remote Sensing Images CLS 1D BD H/V Raster + Shuffle Global GRSL24
@@
HMNet (Xu et al., 2024b) Natural Images Few-shot SEG 2D BD H/W Raster Global NeurIPS24
@@
ViM-UNet (Archit and Pape, 2024) 2D Medical Images (Cell/Neurite) 2D Medical Images SEG 1D BD H Raster Global MIDL24 Short
Swin-UMamba (Liu et al., 2024g) 2D Medical Images (Endo/Micro/MRI) 2D Medical Image SEG 2D BD H/V Raster Global MICCAI24
VM-UNet-V2 (Zhang et al., 2024h) 2D Medical Images (Skin/Endo) 2D Medical Image SEG 2D BD H/V Raster Global ISBRA24
@@
MM-UNet (Du et al., 2024b) 2D/3D Medical Images (MRI) 2D/3D Medical Image SEG 1D SD H Raster Global Scientific Reports24
@@
Polyp-Mamba (Xu et al., 2024c) 2D Medical Images (Endo) 2D Medical Image SEG 2D BD H/V Raster Global MICCAI24
GGVMamba (Zhou et al., 2024d) 2D Medical Images (CXR) 2D Medical Image SEG 1D/2D BD H/V Raster Global MICCAI24
LKM-UNet (Wang et al., 2024b) 2D/3D Medical Images (CT/MRI) 2D/3D Medical Image SEG 1D BD H Raster Global/Local MICCAI24
SegMamba (Xing et al., 2024) 3D Medical Images (CT/MRI) 3D Medical Image SEG 1D BD + SD H + D Raster Global MICCAI24
EM-Net (Chang et al., 2024) 3D Medical Images (CT) 3D Medical Image SEG 1D SD C Raster Global MICCAI24
CAMS-Net (Khan et al., 2025) 2D Medical Images (MRI) 2D Medical Image SEG 1D BD H/C Raster Global WACV25
TP-Mamba (Wang et al., 2024g) 3D Medical Images (CT) 3D Medical Image SEG 1D SD H/V/D Raster Global MICCAI24
ShapeMamba-EM (Shi et al., 2024c) 3D Medical Images (EM) 3D Medical Image SEG 1D SD H Raster Global MICCAI24
PathMamba (Fan et al., 2024) 2D Medical Images (Path) 2D Medical Image SEG 1D BD H Raster Global MICCAI24
Semi-Mamba-UNet (Ma and Wang, 2024) 2D Medical Images (MRI) 2D Medical Image SEG 2D BD H/V Raster Global KBS24
@@
RS3Mamba (Ma et al., 2024d) Remote Sensing Images Semantic SEG 2D BD H/V Raster Global GRSL24
@@
ChangeMamba (Chen et al., 2024e) Remote Sensing Images Change DET/Building Damage Assessment 2D BD H/V Raster Global TGRS24
MambaHSI (Li et al., 2024e) Remote Sensing Images Remote Sensing Images CLS 1D SD H + C Raster Global TGRS24